OpenStudy (anonymous):

Jane surveys a random sample of 200 students from her middle school to learn what type of school lunch is preferred. The table below shows the result of the survey: Lunch Type Sandwich Hot Dog Salad Number of students 84 85 31 Which of the following inferences can be made based on the data? A sandwich is the most preferred lunch. A sandwich is two times more popular than a salad. Sandwiches and hot dogs are almost equally preferred. A hot dog is almost two times more popular than a salad.

OpenStudy (michele_laino):

total number of student: 84+85+31=...

OpenStudy (michele_laino):

please complete my computation

OpenStudy (anonymous):

It is 200 right.

OpenStudy (michele_laino):

ok! right! probability for sandwich is: 84/200=...

OpenStudy (anonymous):

0.42

OpenStudy (michele_laino):

yes! or equivalently 42%

OpenStudy (michele_laino):

next: probability for hot dog, is: 85/200=...

OpenStudy (michele_laino):

the probability is given by the subsequent ratio r: r= favorable outcome/possible outcome now we can consider our probabilities as experimental frequencies

OpenStudy (michele_laino):

\[\Large \frac{{85}}{{200}} = ...?\]

OpenStudy (michele_laino):

next: probability for salad, or experimental frequency for salad is: \[\Large \frac{{31}}{{200}} = ...?\] please complete my computation